OpenStudy (anonymous):

Can someone please factor out (x + 3) cubed?

jimthompson5910 (jim_thompson5910):

it's already factored

jimthompson5910 (jim_thompson5910):

did you mean expand?

OpenStudy (anonymous):

Perhaps I mean expand

OpenStudy (anonymous):

The other issue is expanding (x + delta x) cubed?

jimthompson5910 (jim_thompson5910):

Ok first off we need to expand out (x + 3)(x + 3) (x + 3)(x + 3) = x(x+3) + 3(x+3) (x + 3)(x + 3) = x^2 + 3x + 3x + 9 (x + 3)(x + 3) = x^2 + 6x + 9

jimthompson5910 (jim_thompson5910):

then we use this to expand out (x + 3)(x + 3)(x + 3)

jimthompson5910 (jim_thompson5910):

(x + 3)^3 = (x + 3)(x + 3)(x + 3) (x + 3)^3 = (x + 3)(x^2 + 6x + 9) (x + 3)^3 = x(x^2 + 6x + 9) + 3(x^2 + 6x + 9) (x + 3)^3 = x^3 + 6x^2 + 9x + 3x^2 + 18x + 27 (x + 3)^3 = x^3 + 9x^2 + 27x + 27

jimthompson5910 (jim_thompson5910):

you expand (x + delta x) cubed the same way as well

OpenStudy (anonymous):

I seem to be making an algebraic error in expanding (x + delta x) cubed. Could you please work it out for me. Thank you so much.

jimthompson5910 (jim_thompson5910):

replace delta x with another variable (say y) and try it again

OpenStudy (anonymous):

\[(a+b)³=a^3+3a²b+3ab²+b³\] Just change a to x and b to delta x.

OpenStudy (anonymous):

Well, the problem you see is how Delta X is treated in a multiplicative relation with plain X or X squared.

jimthompson5910 (jim_thompson5910):

not sure what you mean rafirank

OpenStudy (anonymous):

Jim--thank you and thank you very much to Zairhenrique. I've got it now. Thanks to you both. I understand what I have been doing wrong.

jimthompson5910 (jim_thompson5910):

ok glad it's making sense now

OpenStudy (anonymous):

you welcome man